# HCT-QA: A Benchmark for Question Answering on Human-Centric Tables

### v2

@@ 4.2 Analysis and findings
(RQ5) How do HCT properties affect model performance? Table 4 shows the regression coefficient of various HCT properties against Recall (Properties are detailed in Appendix A). We can see that the presence of most of these properties, except Implicit Column Aggregation Terms, decreases the Recall on average. We also see a clear trend in models being able to better understand symmetrical and balanced nesting patterns compared to asymmetrical and unbalanced ones: Symmetric Row Nesting decreases Recall by 10% whereas Asymmetric Row Nesting decreases it by 14%. The same trend is seen when comparing Balanced Column Nesting with Unbalanced Column Nesting with the former having a decrease of 27% on Recall against 31.5% for the latter. Column and row nesting increases the HCT difficulty more than any type of aggregation, having the largest negative impact on Recall. Regression analysis on the full set of properties is given in Appendix D.2.
5 Limitations
The real-world HCTs were collected through various sources, but we could not guarantee a priori a uniform coverage across all layout properties (Section 3.1). However, we controlled the diversity and coverage of the natural language questions using predefined templates to guide their generation. These properties allowed us to characterize specific limitations of the evaluated models (Section 4). The synthetic HCT and QA generators (Section 3.3) allowed scaling up the size of the benchmark datasets at the cost of a lower diversity of the questions and layout structures compared to real-world data. Still, our fine-tuning experiments demonstrate the benefits of the synthetic datasets to improve LLM performances on real HCT-QA test data (Section 4).
6 Conclusion and future work
In this paper, we introduced a benchmark for querying human-centric tables found in PDFs, web pages, and similar documents. Our dataset includes a rich set of real-world and synthetically generated HCTs in addition to thousands of QA pairs, providing a rich resource to evaluate HCTs question-answering methods. Our results show that larger, closed-weight models (e.g., gpt-o1-mini) perform best overall, but struggle with complex queries—especially those involving HCTs with nested structures and queries that require aggregations. Performance improves with increased model size, instruction tuning, and more expressive HTML input formats. Notably, VLMs show promise by effectively processing HCT images. Still, the best LLM (gpt-o1-mini) tops at 69.1% precision and 68.5% recall, while the best VLM (gpt-4o-vision) tops at 48.7% and 62.3%, respectively, letting ample room for improvement. Our preliminary fine-tuning experiments show that fine-tuning Llama-3-8B and FanarC-9B using real-world and synthetic data more than triples their performance compared to no fine-tuning, demonstrating the potential of this approach and the benefit of our synthetic dataset.
Several research avenues could be explored. First, using our HCT-QA benchmark datasets to pre-train or fine-tune foundation models is an important endeavor to enhance their capacities to process HCTs. Our fine-tuning experiments show very promising results in this direction. Second, our work opens the way to address more complex QA tasks on multiple related HCTs spread within and across PDFs and other documents in a data lake. Finally, our synthetic generator could be used for the quantitative evaluation of NL-to-SQL (liu2024surveynl2sqllargelanguage, ), table transformation (Abedjan2016DataXFormerAR_46, ; AutoTables_57, ), and table extraction (Smock2021PubTables1MTC_15, ; tabnet-paper, ) techniques targeting HCTs.
APPENDIX
@@ Appendix A HCT properties
A.1 HCT Property definitions and examples
Table 5 presents some summary statistics about the HCTs in our dataset. Definitions of HCT specific table properties are given in Table 6 and expanded upon with examples in this section.
Table 5: Summary statistics of the proposed benchmark datasets
QNPC Pak. Census US Census ArXiv SYNTHETIC
Average No. Rows 16 20 40 12 11
Average No. Columns 9 7 10 8 14
Average No. Cells 145 146 437 102 151
Average Table Word Count 201 187 575 138 153
Total Number of Tables 958 128 219 575 4,679
Total Number of QA Pairs 7,128 979 387 1,341 67,747
SIGNIFICANT TABLE PROPERTIES FREQUENCY (%)
Standard Relational Table 2.3% 7.0% 6.0% 18.6% 0%
Symmetric Row Nesting 7.4% 0.8% 0.0% 16.2% 100%
Asymmetric Row Nesting 3.2% 0.0% 0.4% 12.0% 0%
Balanced Multi Level Column 60.0% 66.4% 43.0% 59.1% 100%
Unbalanced Multi Level Column 0.8% 3.1% 22.6% 0.8% 0%
Implicit Column Aggregation Terms 6.8% 3.1% 2.6% 0.8% 0%
Implicit Row Aggregation Terms 10.1% 8.6% 43.0% 1.5% 0%
Global Column Aggregation 54.6% 13.3% 4.7% 3.1% 47%
Row Group Label 6.4% 36.7% 35.7% 8.0% 0%
*The percentages represent the occurrence of each table property across datasets.
*This subset of table properties has statistically significant effect on model performance.
Table 6: Definitions of HCT (table) properties.
Property Name Definition
Column Nesting When headers/column names are spread across more than one level (row).
Balanced Column Nesting When the depth of each branch within the nesting is the same throughout the table.
Unbalanced Column Nesting When the depth of each branch within the nesting is not the same.
Symmetric Column Nesting When the child column names are repeated across the table uniformly in columns where nesting is present.
Asymmetric Column Nesting When the child column names are different across the table in columns where nesting is present.
Row Nesting When a cell spans multiple rows but is displayed as one merged cell.
Balanced Row Nesting When the depth of each row branch is the same.
Unbalanced Row Nesting When the depth of each row branch is not the same.
Symmetric Row Nesting When the child cell values are repeated across the table uniformly in rows where nesting is present.
Asymmetric Row Nesting When the child cell values are different across the table in rows where nesting is present.
Column Aggregation Columns that aggregate all or a subset of other columns
Global Column Aggregation A column that aggregates all other numerical columns in the table (excluding local-group aggregation columns).
Local Column-Group Aggregation A column that aggregates a subset of numerical columns in the table.
Explicit Column Aggregation Terms An aggregation column where the column name clearly states it’s an aggregation (e.g., total, sum, percentage).
Implicit Column Aggregation Terms An aggregation column where the column name is ambiguous about whether it’s an aggregation.
Row Aggregation Rows that aggregate all or a subset of other rows
Global Row Aggregation A row that aggregates all other numerical rows in the table (excluding local-group aggregation rows).
Local Row-Group Aggregation A row that aggregates a subset of numerical rows in the table.
Explicit Row Aggregation Terms A row where some textual cell value clearly states it’s an aggregation (e.g., total, sum, percentage).
Implicit Row Aggregation Terms A row where all textual cell values are ambiguous about whether it’s an aggregation.
Standard Relational Table Table displays none of the other HCT properties (nesting, aggregations, labels).
@@ A.1 HCT Property definitions and examples
A.2 Table Annotation Tool
@@
A.3 Understanding the distribution of table properties
@@
Appendix B Question properties
B.1 Query Properties Definitions
Definitions of our query properties are in Table 7.
Table 7: Definitions of query-related properties used in question analysis.
Property Name Definition
Row Filter (RF) Whether the query filters rows based on some condition.
RF Condition Type Lookup The condition filters based on exact matches to cell values.
RF Condition Type Expression The condition uses numeric or logical expressions (e.g., >>, <<).
RF Involved Columns Single Only one column is involved in the filtering condition.
RF Involved Columns Multiple More than one column is involved in filtering.
RF Max Depth Of Involved Columns Maximum header depth of columns involved in filtering.
RF Retained Rows Single Only one row satisfies the filtering condition.
RF Retained Rows Multiple More than one row satisfies the filtering condition.
RF Num Of Conditions Number of conditions used to filter rows.
Returned Columns (RC) Whether the query returns specific columns.
RC Project On Plain Columns are returned without applying expressions.
RC Project On Expression Columns are returned with computed expressions.
RC Max Depth Of Involved Columns Maximum header depth among returned columns.
RC Expression In Table Present Returned expression is already present within the table.
RC Expression In Table Not Present Expression used in the result does not directly appear in the table.
RC Num Of Output Columns Number of columns returned in the query answer.
Yes/No Whether the question expects a yes/no answer.
Yes/No Scope ’Single’ if the yes/no judgment is based on one row or column, ’Multiple’ if it depends on multiple rows/columns.
Aggregation Whether the query includes any aggregation operation.
Aggregation On Plain Aggregation is applied on column values as they are.
Aggregation On Expression Aggregation is applied on computed expressions.
Aggregation Details Specifies the type of aggregation (e.g., min, max, count, sum, average).
Aggregation Grouping Local Aggregation is grouped by some column values.
Aggregation Grouping Global Aggregation is done over all table rows (no grouping).
Aggregation In Table Present The aggregated result already exists within the table.
Aggregation In Table Not Present The aggregation result is not directly present in the table.
Aggregation Num Of Aggregations Number of different aggregations needed to answer the query.
Rank Whether the question involves ranking or ordering.
Rank On ’Plain’ if ranking is performed directly on a column, else ’Expression’ if ranking is based on computed values or expressions
B.2 Human Verification Process
To ensure the accuracy and clarity of model-generated queries, we implemented a rigorous human verification and correction process. This process was designed to both evaluate the quality of the generated queries and to produce a high-quality dataset of verified query-answer pairs.
B.2.1 Annotation Protocol
Each model-generated query was evaluated by two independent annotators. Annotators were presented with the query, the associated table, and the model-provided answer. They were asked to complete two tasks:
  1. (i)
    Query Quality Rating: Annotators rated the query as one of the following:
    • Good: The query was well-formed, unambiguous, and meaningful with respect to the table.
    • Ambiguous: The query could be interpreted in multiple ways or was under-specified.
    • Nonsensical: The query did not make logical sense given the table content or was completely unrelated.
  2. (ii)
    Answer Correction: Annotators were instructed to check if the provided answer matched the correct answer, based solely on the table. If the answer was incorrect, they provided the corrected version.
If annotators disagreed on the quality label or the answer correction, a third annotator adjudicated the final decision.
B.2.2 Annotator Quality Control
To validate the consistency and attentiveness of annotators, we included a set of human-generated queries whose correctness was manually verified beforehand. Annotators who incorrectly marked more than one of these as ”nonsensical” or flagged correct answers as incorrect were removed from the pool. This vetting procedure ensured a reliable annotation process.
B.2.3 Query Outcome Statistics
After annotation:
  • 4% of model-generated queries were marked as nonsensical.
  • 24% were marked as ambiguous and subsequently rewritten.
  • 72% were rated as good and used as-is.
  • Approximately 50% of the model-generated queries required answer correction.
B.2.4 Examples of Query Types
Good Query:
What was the total number of passengers in 2019?
(Clearly refers to a valid column and year in the table; unambiguous and directly answerable)
Ambiguous Query:
How many were there in 2020?
(Unclear what ”many” refers to—passengers, flights, delays? Multiple possible interpretations)
Rewritten Version of Ambiguous Query:
How many total passengers were recorded in 2020?
Nonsensical Query:
what was the year for passengers and plane and its year ?
(Combines unrelated concepts not present in the table; not meaningful)
B.3 Query Generation and Annotation Query Tool
Figure 25: A user-friendly tool for query generation and their properties.
We developed a tool for QA generation shown in Figure 25. The left side displays the HCT table, while the right side enables the annotator to create a question and identify its properties. As shown in the figure, we capture the “Row Filter” property, which corresponds to the WHERE Clause in the SQL statement, where we capture the condition type (“Lookup” or “Expression”), the number of columns involved in the conditions, if the table involves multi-level column names, we capture the max level depth in the conditions, and the number of retained rows after all conditions are applied. The “Returned Columns” property corresponds to the SELECT clause in the SQL statement, where we capture the number of columns produced in the output, whether or not expressions are involved, and the max level depth among the output columns.
More complex properties involve “Aggregation”, “Rank”, and “Yes/No”, where we capture more details for each property as illustrated in Figure 25. For example, for “Aggregation”, we capture whether it is global or local, the number of involved aggregations, and the type of the aggregation. On the other hand, for the “Yes/No” query, which could involve many of the other properties (e.g., filtering, aggregation, and rank), we additionally capture whether the final Boolean expression is evaluated on a “Single” row or across “multiple” rows.
B.4 Understanding the distribution of query properties
The Figure 26 and Figure 27 show the distribution of the top-$5$ main and detailed real question properties per dataset.
The main question property distributions in Figure 26 are quite similar across datasets, with similar proportions of Returned.Columns, Row.Filter, and Aggregation properties. The proportion of the Returned.Columns and Row.Filter combination is more prominent in QNPC and Pak. C than in the other two datasets.
Regarding the detailed question properties in Figure 27, the distribution of top-$5$ properties is slightly decreasing for all datasets, while there are a few but frequent complex patterns in all datasets, and the remaining patterns are rare.
Refer to caption Refer to caption
(a) QNPC (b) Pak. C
Refer to caption Refer to caption
(c) US Census. (d) ArXiv
Figure 26: Distribution of main question properties for each real dataset.
Refer to caption Refer to caption
(a) QNPC (b) Pak. C
Refer to caption Refer to caption
(c) US Census. (d) ArXiv
Figure 27: Distribution of detailed question properties for each real dataset.
Appendix C Synthetic Data Generation
Refer to caption Refer to caption
(a) (b)
Figure 28: Overview of synthetic tables generation. (a) We generated $4,679$ unique synthetic HCTs from 77 semantics with up to $15$ questions each, leading to $67,747$ benchmark data triplets $(T_{HCT},Q_{NL},A)$. (b) These triplets derive from ground-truth data triplets $(T_{REL},Q_{SQL},A)$: Tables $T_{HCT}$ are obtained by pivoting relational tables $T_{REL}$; Natural language questions $Q_{NL}$ are obtained from templates of SQL queries $Q_{SQL}$; and the ground-truth answer AA is obtained by running $Q_{SQL}$ on $T_{REL}$.
Figure 29: Details of synthetic tables generation. User-defined semantic data, table and question templates are used to generate instances of pairs $(T_{REL},T_{HCT})$ of relational tables and pivoted human-centric variants as well as corresponding pairs $(Q_{SQL},Q_{NL})$ of templated SQL queries and NL questions. All the values (dashed color frames) of columns A, B, C, and D of the relational table are factorized as row and column headers in the HCTs (HCT1, HCT2). In contrast, the whole value column E (blue frame) is reproduced in the core area of the HCT (HCT3). In this instance, the ground-truth answer $A$ to the question "What is the amount of Export of Milk in 2017?" ($Q_{NL}$) related to $T_{HCT}$ is "57" (Dashed blue frame). This answer AA is obtained by running the SQL query $Q_{SQL}$ paired with $Q_{NL}$ onto $T_{REL}$. See Section C.1 for more details.
The statistics of the semantics of the synthetic tables and questions we generated are listed in Figure 28a. A summary view of the synthetic generation process is displayed in Figure 28b.
In this section, we explain the synthetic table and QA generation process in Section C.1. Then we detail the configuration templates for semantics (Section C.2), tables (Section C.3), and questions (Section C.4).
C.1 Generating process
The first step of the process is to manually define three different sets of templates (Symbols refer to Figure 29):
Semantic data
are textual items used as attributes and values to form realistic relational tables (e.g., ”Year”: ”2015”, ”2016”, ”2017”; ”Food category”: ”Beverage”, ”Dairy”, ”Cereals”;…). Details of this data structure are given in Section C.2.
Table templates
group several of the semantic data to form a family of pairs ($T_{HCT},T_{REL}$) of HCTs and corresponding relational tables. For instance, to generate a ”Food import-export” HCT, we first indicate the use of semantic data ”Year”, ”Import-Export”, food ”Category” and ”Item” (Columns A, B, C, D of $T_{REL}$, HCT1 and HCT2 headers in $T_{HCT}$), and then specify the number and complexity of the tables to generate in terms of (i) the range of numerical values to add in the core of the table (HCT3, ”Value” column E of $T_{REL}$), (ii) the depth of the column and row headers (HCT1, HCT2) and values therein, (iii) the presence and type of aggregation in $T_{HCT}$ (”Average” in HCT1 and HCT2) and (iv) other visual aspects (e.g., indented headers). Several format parameter values (ii-iv) are randomly generated from the same relational table instance $T_{REL}$ to vary the appearance of the HCTs without changing its content. Details of these table templates are given in Section C.3.
NL++SQL templates
are natural language question templates paired with SQL query templates, both related to a specific table template. There are $15$ different templates with increasing levels of complexity. For instance, the template:
What is the amount of_\$Import_Export of__\$Item__\$Category in__\$Quarter==of==\$Year? will generate the following natural language question based on the items selected from the current table instance:
What is the amount of Export of Milk in 2017?” that will be used as a prompt for the evaluated LLM along with the $T_{HCT}$ instance. The corresponding SQL query is:
$Q_{SQL}$: "SELECT Value FROM DBdata WHERE((Import_Export = ’Export’ AND Year = ’2017’))
AND ((Category = ’Dairy’ AND Item = ’Milk’))"
and will be run on the $T_{REL}$ instance paired with $T_{HCT}$ to get the ground-truth answer A=57A="57".
Details of these question templates are given in Appendix C.4.
Afterward, referring to Figure 29, we automatically process these template files to generate a large number of HCT and question instances: table templates use (A) semantic data to generate table instances (B) varying randomly or systematically the template parameters to specify the semantic data (C) to be used in rows (HCT1) or columns (HCT2), the aggregates, and the values (HCT3) of that table instance. From each table instance, we generate a relational table $T_{REL}$~(D) and one of its randomized HCT representations $T_{HCT}$ (E) using the pivottabler [9] R package. We also use its related set of question templates (F) to generate both SQL queries $Q_{SQL}$ and their natural language templated versions $Q_{NL}$ (G). We run the SQL query on the relational table instance $T_{REL}$ (H) to get the ground-truth answer $A$ (I), valid by design for the natural language question. There are $15$ question templates for each table template, leading to up to $15$ types of questions for each HCT. The generated table values (Column E ”Value” in $T_{REL}$, HCT3 in $T_{HCT}$) have no duplicate to avoid ambiguities when evaluating the matching between the ground-truth and LLM-predicted answers. Some semantic data are hierarchical, such as food categories and items (HCT2, columns C,D); we ensure they cannot be split across row and column headers of the HCT.
Here are the $15$ types of natural language questions $Q_{NL}$ generated for the table $T_{HCT}$ in Figure 29 based on templates (”,” and ”;” in answers, are column and row separators respectively):

### v3

@@ V-B Analysis and findings
(RQ5) How do HCT properties affect model performance? Table VII shows the regression coefficient of various HCT properties against F1 scores. We can see that the presence of most of these properties, except Implicit Column Aggregation Terms, decreases the F1 score on average. We also see a clear trend in models being able to better understand symmetrical and balanced nesting patterns compared to asymmetrical and unbalanced ones: Symmetric Row Nesting decreases F1 by 10 percentage points whereas Asymmetric Row Nesting decreases it by 14 percentage points. The same trend is seen when comparing Balanced Column Nesting with Unbalanced Column Nesting with the former having a decrease of 27 percentage points on F1 against 31.5 percentage points for the latter. Column and row nesting increases the HCT difficulty more than any type of aggregate rows or columns, having the largest negative impact on F1 scores.
(RQ6) How do question properties affect model performance? Our question metadata allows for fine-grained analysis to understand what types of questions different models excel at and struggle with. Figure 9 shows such an analysis on a subset of models and metadata. Heatmap (A) shows that all models struggle with questions requiring aggregations or expressions on columns. We can also see that Qwen3-14B, despite being better than models like Fanar-C-9B and gemma-3-27b at questions involving ranking and row filtering, pales in comparison to both on yes/no questions. Furthermore, Qwen2.5-72B shows good performance on aggregation questions compared to other models. In general, models perform better with yes/no and row filter questions as this type of question is answered directly from the data cells, similarly to scan, filter and project stateless operators in databases (i.e., a single row/cell is enough to answer the question), while ranking is more challenging for the models because, similar to the ORDER BY stateful operator in databases, it requires a full view of the data. Aggregation is even more challenging because it operates on multiple cells, which can be in different locations of the HCT. In Figure 9), we observe that, for example, end users would be better off using gemma-3-27b than the larger Llama-3.3-70B and Llama-3.1-70B models for questions requiring aggregation, row filtering, and yes/no answers, as their performance is nearly the same, while gemma-3-27b is much more efficient due to its smaller size.
We go one step deeper and analyze what types of aggregations these models struggle with in particular. Heatmap (B) in Figure 9 shows the F1 scores of the models on the five different types of aggregation. We can clearly see that all models struggle significantly with averaging, presumably because this involves a multi-step calculation. We can also see that both Llama-3.1-70B and Llama-3.3-70B are particularly good at min and max questions but struggle more with sum questions. Whereas, Qwen2.5-72B is the only model that performs reasonably well on sum questions which would explain how it outperforms the other models on aggregation questions in general. As expected, all models struggle with questions involving aggregation and particularly those that require averaging values. The metadata provided in HCT-QA allows for such granular analysis to be done for each individual model so that specific areas for improvement can be identified.
(RQ7) How are model size and performance correlated? Figure 10 shows how F1 score changes between different size variants of the same model. As expected, larger models consistently outperform their smaller counterparts. The gemma-2 models have the sharpest increase with $0.40\%$ F1 per 1B parameters (F1//B). Even when scaling to larger sizes like with Qwen2.5 and Llama-3.1 we see an increase of $0.34\%$ F1//B and $0.37\%$ F1//B , respectively. Hence, it is worth exploring fine-tuning smaller models for this task rather than simply relying on larger, resource-hungry off-the-shelf models.
Figure 10: F1 score across model size (parameters) within model variants
VI Conclusion and future work
In this paper, we presented HCT-QA, a comprehensive benchmark designed to evaluate language models on question answering over human-centric tables (HCTs), typically embedded within documents and web pages.
Our benchmark reveals several key insights. First, both LLMs and VLMs exhibit strong potential for question answering on HCTs. Second, finetuning LLMs substantially enhances performance, even when finetuned on only synthetic data, models demonstrated strong generalization to real-world HCTs. Finally, even the best-performing models (e.g., ChatGPT 4o) continue to face challenges with complex queries (e.g., involving aggregations) on complex HCT structures. Futhermore, our benchmark can be beneficial to other researchers in at least two ways: (i) Using the synthetic data generator to create training and test HCTs and QA pairs for new thematic domains in a cheap and quick manner. This is especially useful for training, finetuning, and evaluating new models. (ii) Adding question templates to the generator to cover more types of questions.
Future Work. Our benchmark can be extended along several complementary directions. First, we plan to broaden the scope of supported queries by enabling cross-HCT question answering, e.g., queries involving joins or unions across multiple HCTs. We also aim to incorporate queries that require more advanced OLAP operations, including CUBE and PIVOT. Initial experiments with two medium-sized LLMs indicate that these models are currently unable to handle such operators, highlighting substantial room for improvement.
Second, we intend to conduct deeper studies on fine-tuning strategies for both LLMs and VLMs to improve their structural understanding of HCTs. In addition, we will evaluate emerging systems that go beyond current LLM/VLM architectures, as these models may offer improved reasoning capabilities for complex tabular and visual data.
Third, we plan to enhance the linguistic diversity of the synthetic NL questions. One direction is use question paraphrasing capabilities using LLMs, which would increase lexical variation in the generated QA pairs. This extension will require robust validation strategies to ensure semantic equivalence between the original and paraphrased questions. Furthermore, the generator can be enriched with additional NL templates that capture implicit comparative terms such as most, least, highest, and lowest, instead of relying solely on explicit SQL-style operators such as MIN and MAX.
VII AI-Generated Content Acknowledgment
Here are the different places where we used AI models in our work:
  • LLMS were used to generate natural language questions and answers for the benchmark dataset as explained in Sections III-A.
  • The LandingLens from LandingAI tool was used for automated extraction and conversion of tabular data from PDF documents into HTML format (Section III-A).
  • Azure AI Document Intelligence was used to extract table images from PDF documents (Section III-A).
@@ Appendix A HCT properties
A-A HCT Property definitions and examples
Definitions of HCT specific table properties are given in Table II and expanded upon with examples in this section.
@@ A-A HCT Property definitions and examples
Figure 13: Example of Unbalanced Column Nesting
Figure 14: Example of Balanced Column Nesting
@@
A-B Table Annotation Tool
@@
A-C Understanding the distribution of table properties
@@
Appendix B Synthetic Data Generation
B-A Synthetic Generator
In the following subsections, we detail the configuration templates we used for setting the domain vocabulary (Section B-B) common to both HCTs, relational tables, SQL queries, and natural language questions. Then we detail the templates used for generating the synthetic HCTs (Section B-C), and those used for generating the SQL queries and the Natural Language questions (Section B-D).
The codes of the synthetic generator and the instructions for installing and running them are available at https://github.com/qcri/HCTQA-Benchmark/blob/main/synthetic_data_generator/README_SYNTHETIC_GENERATOR.md.
B-B Domain vocabulary templates
Figure 33: Domain vocabulary for generating synthetic tables and questions. Examples of $T_{REL}$ (a) and a corresponding $T_{HCT}$ (b) related to the amount of food import-export in some place. (c) JSON format to describe Food Import-Export domain vocabulary used to generate these tables (see Section B-C) and the related questions (see Section B-D). Not all vocabulary is used in $T_{REL}$ and $T_{HCT}$.
In this section, all terms refer to Figure 33.
Our synthetic data relies on domain vocabulary naming the dataset’s attributes and values. A synthetic $T_{HCT}$ (Figure 33b) is obtained by pivoting a relational table $T_{REL}$ (Figure 33a) composed of at least two nominal attributes and a single numerical attribute always named Value.
There are two types of nominal attributes:
  • Hierarchical attributes like food Category (e.g. Meat, Cereals, Dairy, Beverage) and food Item (e.g., Cream, Soda, Milk) are dependent on each other. If both attributes appear in the HCT, the food Item must be displayed at a lower level than the food Category, i.e. closer to the inner part of the HCT, and all must appear together either in row headers or column headers.
Domain vocabulary is encoded in a JSON file (Figure 33c) with two main keys: data and values.
  • data refers to a list of attributes (e.g., ImportExport, Years, Quarters) or groups of attributes (e.g., Food), each described with three keys:
    • code is the primary key used in table and question templates to identify the independent attributes (e.g., ImportExport or Years) and the hierarchical groups of attributes (e.g., Food).
    • names is the list of attribute names displayed in $T_{REL}$ column headers but typically hidden (implicit) in $T_{HCT}$ and $Q_{NL}$. It is a list with a single name for independent attributes, or with as many names as levels for hierarchical groups of attributes (e.g., Food), ordered from top (outer) to bottom (inner) levels (e.g., Category, Item).
    • values is the list of values (e.g., 2015, 2016... for the attribute Year). It can also be the hierarchy of values of a group of attributes. For instance, in Food group, Meat is a value of the attribute Category, and Beef, Chicken…, are values of the attribute Item, whose parents are Meat in the Food hierarchy. Attribute values are always displayed as row or column headers in $T_{HCT}$.
  • values refers to the numerical values to be generated in both the core part of $T_{HCT}$ and the Value column of $T_{REL}$. This object contains shortcut keys (e.g., realUnit or intSigned) to generate specific ranges and types of numbers when called in table templates. The value associated with such shortcuts must be a list of two numbers $[m,M]$ representing the minimum $m$ and maximum $M$ possible values. If any of these min and max is a decimal number, the generated value will be a real number with two decimal digits drawn uniformly in this interval. If both min and max are integers, the generated value will be an integer drawn uniformly in this interval. The generator will ensure no replica in the generated values, reals or integers, possibly overflowing the max value if asked to generate more numeric values than the range allows.
Customization of domain vocabulary All terms in domain vocabulary are customizable except the JSON keys data, code, names, and values, following the rules above. The number of decimals of generated real numbers is specified by NUM_DECIMAL_DIGITS_REAL_FORMAT in config.R. The domain vocabulary is stored in the file PARAM_semantics.json.
B-C Table templates
The generic table templates described in the file PARAM_tableTemplate.json are used to generate a (large) set of $T_{REL}$ and $T_{HCT}$ tables from the domain vocabulary (See Section B-B).
This template contains the following keys:
  • tables is a list of table content specifications (values, headers, and aggregates) with coherent domain vocabulary. Each initial table template in the list has the following keys:
    • values is either a shortcut name referring to one of the keys (e.g., realUnit) stored in values in the domain vocabulary (see Figure 33), or directly an interval of the form $[m,M]$ following the description of the main key values in Section B-B
    • valueUnit (Place-holder for a future version, not yet implemented) is a character string to be added to give units to values in the core part of the table (e.g., $\%$ sign after each value in the core part of $T_{REL}$).
    • rowCodes is a list of one or more of the data code keys (e.g., Food or Quarters) listed in the domain vocabulary. The generator will pick among these codes to generate the rows of $T_{HCT}$, and one part of the nominal attributes of $T_{REL}$.
    • rowSamples is a list containing as many intervals $[m,M]$ as there are codes listed in rowCodes, in the same order. Each interval specifies the minimum $m$ and maximum $M$ number $n$ of values to select among the list $L$ of values of the specified attribute. The generator first picks $n$ at random in $[m,M]$, then selects $n$ consecutive values in $L$ starting at a random position in $L$ but ensuring there are always $n$ values in the result if the list $L$ is long enough, or return $L$ itself otherwise. For instance, let’s consider an independent attribute Years with a list of values: $L=[2001,2002,2003,2004,2005,2006]$, if we set rowSamples to $[1,5]$, and the generator randomly picks $n=3$, it then picks a random index $i$ between $1$ and $4$, e.g., $i=2$, to generate the final subset $[2002,2003,2004]$ of years to use in the table. For hierarchical attributes, the same process applies recursively to all levels of the hierarchy, starting from the top level. If the interval is set to $[0]$ instead, all the values of the corresponding independent attribute are generated, or all the values at all hierarchy levels are generated for hierarchical attributes. Note that the col_row_levels parameter limits the depth of row and column headers, so the attributes finally displayed in $T_{HCT}$.
    • colSamples is the same as rowSamples but for columns.
    • agg_name1 is the name of the aggregate to display in the table.
These generic table templates PARAM_tableTemplate.json are used with the domain vocabulary PARAM_semantics.json to generate individual templates using the script 1_geneAllTablesJSONfromPatterns.R. The individual templates generated and described below are stored in the file PARAM_tablesToGenerate.json.
Customization of generic table templates These generic table templates are customizable except for all the JSON keys and following the rules above.
An individual table template is a JSON object instantiating one combination of the parameters shuffle, col_row_levels, col_row_name_pos, col_row_agg_pos, and row_format common to all initial table templates, and set in the generic table template. It contains the following keys:
  • replica is identical to replica set in the generic table template.
  • agg_fun1 is identical to agg_fun1 set in the generic table template.
  • values is identical to values set in the generic table template.
Two examples of individual templates with two of the five sampled $T_{REL}$ and corresponding $T_{HCT}$ are shown in Figure 34.
Customization of individual table templates There is no possible customization of these individual table templates automatically generated from the generic table templates.
B-D Natural Language question templates
We defined $15$ SQL templates in Table III and transcribes them into the HCT domain using $15$ domain-specific Natural Language templates (Figures 36 and 37 show all the natural language question templates in JSON format related to the seven generated HCT domains).
Figure 35: Example of a food import-export domain $T_{REL}$ (left) and $T_{HCT}$(right) to explain $Q_{NL}$ templates. All the values (dashed color frames) of columns A, B, C, and D of the relational table are factorized as row and column headers in the HCTs (HCT1, HCT2). In contrast, the whole value column E (blue frame) is reproduced in the core area of the HCT (HCT3). In this instance, the ground-truth answer AA to the question "What is the amount of Export of Milk in 2017?" ($Q_{NL}$) related to $T_{HCT}$ is "57" (Dashed blue frame). This answer AA is obtained by running the SQL query $Q_{SQL}$ paired with $Q_{NL}$ onto $T_{REL}$.
For instance, the $Q_{NL}1$ template below transcribes the $SQL_{1}$ template (Table III) for the Food Import-Export domain HCT displayed in Figure 35c:
What is the amount of_ \$Import_Export of__ \$Item__ \$Category in__ \$Quarter==of==\$Year?
This template is instanciated with detailed row and column attribute values of the HCT, to generate the following natural language question:
What is the amount of Export of Milk in 2017?
Notice that not all the table attributes may be displayed in the HCT, and the values vary due to random selection of the HCT instance.
This question will be used as a prompt for the evaluated LLM along with the $T_{HCT}$ instance. The corresponding SQL query instance is:
SELECT Value FROM DBdata
WHERE ((Import_Export = ’Export’ AND Year = ’2017’))
AND ((Category = ’Dairy’ AND Item = ’Milk’));
and it will be run on the $T_{REL}$ instance paired with $T_{HCT}$ to get the ground-truth answer A=57A="57".
Here are the $15$ types of natural language questions $Q_{NL}$ generated for the table $T_{HCT}$ in Figure 35 instantiated from the template of Figure 36a (”,” and ”;” in answers, are column and row separators respectively):